You are buying a car. You do not like playing the financing game where they leave you in the sales room for an hour to go calculate everything – because you are busy, you got stuff to do. The car costs $15,000. The interest on the car is 2.5% annual. You want to pay it off in 5 years. Tell the salesperson what the car payment will be per month.

Answer: $180.45

Answer: $266.21

Answer: $789.38

Answer: $50.00
Answer #1

This question requires application of PV of annuity, which is mathematically represented as:

For this question, n = 5 * 12 = 60 months; r = 2.5%/12 = 0.2083% (monthly)

15,000 = P * [\frac{1 -(1+0.002083)^{-60}}{0.002083}]

15000 = P * (0.1174/0.002083)

P = 15000/56.3464 = $266.21 (Answer)